In a purely inductive circuit, the rms voltage is
Answer #1

here Irms=4.32A Vrms=102v F=57.6

w=2*3.14*F

w=2*3.14*57.6

w=361.728

Im=Vm/wL

1.41*Irms=1.41Vrms/361.728*L

L=0.065273 H Ans